OpenStudy (anonymous):

Use a clock or stopwatch to take your pulse rate while sitting relaxed. Record the number of beats for thirty seconds. Now, walk out your front door and run around your house, then come back inside. Take your pulse again and record the number of beats for thirty seconds. What is the percent increase in your pulse rate? Show your work. Round your percent to the nearest whole number.

OpenStudy (anonymous):

when I was relaxed my pulse was 24 beats in 30 seconds after I ran around my house my pulse was 57 beats in 30 seconds

OpenStudy (anonymous):

plzzzzz will give medal

OpenStudy (mathstudent55):

\(percent~change = \dfrac{new~number - old~number}{old~number} \times 100\) Use 57 for the new number and 30 for the old number.

OpenStudy (anonymous):

I don't understand

OpenStudy (mathstudent55):

The right side of the equal sign is a fraction multiplied by 100, right?

OpenStudy (mathstudent55):

The heart rate was 24 and then became 57. The "old number" is 24. The "new number" is 57. Where you see "old number" in the fraction, replace it with 24. Where you see "new number" in the fraction, replace it with 57. Then do the subtraction in the numerator. Then divide the numerator by the denominator. Then multiply the result by 100. The number you have now is the percent change. Since it is positive, it is a percent increase.

OpenStudy (mathstudent55):

Example: I had 85 stamps in my collection. After I received a birthday present of stamps, now I have 102 stamps. What is the percent increase in the number of stamps from the birthday present? Here, the new number is 102, and the old number is 85. \(percent~change = \dfrac{new~number - old~number}{old~number} \times 100\) \(percent~change = \dfrac{102 - 85}{85} \times 100\) \(percent~change = \dfrac{17}{85} \times 100\) \(percent~change = 0.2 \times 100\) \(percent~change = 20\) The percent increase is 20%.

OpenStudy (anonymous):

thank you so much @mathstudent55

OpenStudy (mathstudent55):

You're welcome. What percent increase did you get?

OpenStudy (anonymous):

I did not do it yet to be honest I am still trying to figure out the fraction part

OpenStudy (anonymous):

the old number goes on the top right

OpenStudy (mathstudent55):

I'll do it in a drawing to show you more clearly where everything goes.

OpenStudy (anonymous):

ok thanks

OpenStudy (mathstudent55):

Do you follow everything in the drawing so far? 

OpenStudy (anonymous):

 right